OpenStudy (anonymous):

how do you do 2x(x-3)=5x^2-7x????

OpenStudy (anonymous):

Solve the equation and set equal to 0: 2x(x-3)=5x^2-7x 2x^2-6x=5x^2-7x 0 = 3x^2-x 0 = x(3x-1) So set each part equal to 0, you get x=0 and x=1/3!

OpenStudy (anonymous):

where did the other ^2 come from i been trying to get it and i was no where near that ansa??

OpenStudy (anonymous):

When you distribute the 2x into the quantity (x-3) You're doing: 2x(x) + 2x(-3) = 2x^2 - 6x

OpenStudy (anonymous):

but i still didnt come up with that ansa i was trying to see where i went wrong...and thanks!!

OpenStudy (anonymous):

Not a problem buddy!

OpenStudy (anonymous):

i see i forgot to distribute the final part x(3x-1) smh thanks again